# Cross-Stock Predictability via LLM-Augmented Semantic Networks

### 21 Apr 2026 v1

@@ 1 Introduction
Cross-stock predictability—the phenomenon that a stock’s future return can be partially predicted by the contemporaneous behaviour of economically linked firms—is one of the most robust empirical regularities in asset pricing Cohen and Frazzini (2008); Menzly and Ozbas (2010). This predictability arises because information diffuses gradually across interconnected firms: a demand shock to a downstream manufacturer, for instance, eventually propagates to its upstream suppliers, but often with a delay that creates exploitable price dislocations. Capitalising on this phenomenon requires a reliable map of economic linkages—a network that captures who is genuinely connected to whom.
@@ 2.1 Cross-Stock Predictability and Economic Linkages
A rich literature in finance documents that the returns of economically linked firms can predict the future returns of a focal firm, a phenomenon often attributed to investor inattention and slow information diffusion. Researchers have identified various predefined linkages, including supply chains Cohen and Frazzini (2008), industry classifications Moskowitz and Grinblatt (1999), geographic proximity Parsons et al. (2020), and shared analyst coverage Ali and Hirshleifer (2020). Recent evidence further documents cross-stock predictability based on return similarity (Chen and Fan, 2026).
@@ 2.3 LLMs for Financial Reasoning
Large Language Models (LLMs) have shown remarkable capabilities in financial text analysis, moving beyond simple sentiment analysis to complex reasoning tasks and interpretable internal representations Wu et al. (2023); Kong et al. (2024a, b); Chen et al. (2025a); Xie et al. (2024). Recent applications include using LLMs for direct stock return prediction Chen et al. (2023); Lopez-Lira and Tang (2023), macroeconomic forecasting Chen et al. (2025b), corporate event analysis Xie et al. (2023), extracting structured investor beliefs Gao et al. (2024), and tracking evolving semantic signals in corporate disclosures Choi et al. (2025). Furthermore, the frontier of LLM application is shifting from passive analysis to agentic AI frameworks that autonomously discover and validate systematic trading factors Huang and Fan (2026). However, recent studies raise critical concerns regarding direct economic forecasting, warning that the apparent predictive success of LLMs may largely stem from data contamination and memorization rather than genuine foresight Lopez-Lira et al. (2025).
Our work extends this paradigm and circumvents the memorization problem by deploying LLMs not as direct return predictors, but as semantic reasoning agents for network refinement. Unlike traditional NLP approaches that rely on unweighted embedding similarities, our Retrieve-then-Reason framework aligns with the financial literature’s call for distinguishing the nature of cross-stock relationships Yan and Yu (2023). By prompting the LLM to explicitly classify edges into categories (e.g., Competitor, Supply Chain, Complementary) and filtering out noisy competitor links, we construct a high-fidelity economic graph. This approach directly addresses the limitations of embedding-based networks and mitigates the lookahead bias associated with direct LLM forecasting, providing a structurally sound foundation for cross-stock signal aggregation.

### 26 Apr 2026 v2

@@ 1 Introduction
Cross-stock predictability—the phenomenon that a stock’s future return can be partially predicted by the contemporaneous behaviour of economically linked firms—is one of the most robust empirical regularities in asset pricing (Cohen and Frazzini, 2008; Menzly and Ozbas, 2010). This predictability arises because information diffuses gradually across interconnected firms: a demand shock to a downstream manufacturer, for instance, eventually propagates to its upstream suppliers, but often with a delay that creates exploitable price dislocations. Capitalising on this phenomenon requires a reliable map of economic linkages—a network that captures who is genuinely connected to whom.
@@ 2.1 Cross-Stock Predictability and Economic Linkages
A rich literature in finance documents that the returns of economically linked firms can predict the future returns of a focal firm, a phenomenon often attributed to investor inattention and slow information diffusion. Researchers have identified various predefined linkages, including supply chains (Cohen and Frazzini, 2008), industry classifications (Moskowitz and Grinblatt, 1999), geographic proximity (Parsons et al., 2020), and shared analyst coverage (Ali and Hirshleifer, 2020). Recent evidence further documents cross-stock predictability based on return similarity (Chen and Fan, 2026).
@@ 2.3 LLMs for Financial Reasoning
Large Language Models (LLMs) have shown remarkable capabilities in financial text analysis, moving beyond simple sentiment analysis to complex reasoning tasks and interpretable internal representations (Wu et al., 2023; Kong et al., 2024a, b; Chen et al., 2025a; Xie et al., 2024). Recent applications include using LLMs for direct stock return prediction (Chen et al., 2023; Lopez-Lira and Tang, 2023), macroeconomic forecasting (Chen et al., 2025b), corporate event analysis (Xie et al., 2023), extracting structured investor beliefs (Gao et al., 2024), and tracking evolving semantic signals in corporate disclosures (Choi et al., 2025). Furthermore, the frontier of LLM application is shifting from passive analysis to agentic AI frameworks that autonomously discover and validate systematic trading factors (Huang and Fan, 2026). However, recent studies raise critical concerns regarding direct economic forecasting, warning that the apparent predictive success of LLMs may largely stem from data contamination and memorization rather than genuine foresight (Lopez-Lira et al., 2025).
Our work extends this paradigm and circumvents the memorization problem by deploying LLMs not as direct return predictors, but as semantic reasoning agents for network refinement. Unlike traditional NLP approaches that rely on unweighted embedding similarities, our Retrieve-then-Reason framework aligns with the financial literature’s call for distinguishing the nature of cross-stock relationships (Yan and Yu, 2023). By prompting the LLM to explicitly classify edges into categories (e.g., Competitor, Supply Chain, Complementary) and filtering out noisy competitor links, we construct a high-fidelity economic graph. This approach directly addresses the limitations of embedding-based networks and mitigates the lookahead bias associated with direct LLM forecasting, providing a structurally sound foundation for cross-stock signal aggregation.